Which statement is true?
#1 The equation –3|2x + 1.2| = –1 has no solution.
#2 The equation 3.5|6x – 2| = 3.5 has one solution.
#3 The equation 5|–3.1x + 6.9| = –3.5 has two solutions.
#4 The equation –0.3|3 + 8x| = 0.9 has no solution.

Answer by tmwilliamlin168(27)

1st statement:
-3(2x+1.2)=-1 OR 3(2x+1.2)=-1

-6x-3.6=-1 6x+3.6=-1
1st statement has 2 solutions
2nd statement:
absolute values either have 0 or 2 solutions
3rd statement:
absolute value multiplied by a postive number can't be negative
4th statement:
absolute value times negative number can't be a positive number so it is this